The sum of number,3/4 of the number and 5/8 of the number is 19.find the number
Answer by suruman(21)

let the number be x:
x + (3/4)*x+(5/8)*x = 19
Multiplying 8 by both sides,
8x + 6x + 5x = 152
19x = 152
x = 152/19 = 8
Hence the number is 8.